Minutes of the VIS Meeting on 2016/6/10
2016/6/10 13:30 - 15:00 @ Room 210
Participants:Takahashi, Ishizaki, Sato, Tanaka, Fabian, Mark, Fujii, Hirata, Shoda, Akutsu
Simulation: Ask Okutomi the status (Takahashi)
- Structural analysis of the frame lock was done.
- EQ stop test
Confirm what needs to be done with Ishizaki-san
Hurry up the production of the frame lock (Takahashi)
- Accelerometers check
- Dismounted them from 3 of the TAMA SAS
- Will work on the remaining 1 soon.
- Standard filer
- Measuring the optimal loads
- Bottom Filter
- Calculated the size of the blades
- Finish the cutting by the end of July
- Release sensor
- Waiting for the samples to test
- Type B Past Week Summary
- Fabian finished dummy BS assembly.
Mark and Tatsumi-san glued magnets into both types of BS optic flags, and discs into BS IM flags. klog Two optic flags had to be redone because the magnets moved while the glue was drying. Hirata-san helped remove the magnets with acetone in the ultrasonic cleaner. klog
Tatsumi-san, Fabian and Mark took the dummy BS to Kashiwa and glued the flags and prisms. klog We will check status on Mon 6/13.
- Hirata-san moved in another filter for testing (but didn't get to measure it).
- Hirata-san washed a batch of parts.
- Hirata-san modified the design of the parts of the assembly frame supporting the BF and SF to include bushings for smoother movement.
- Hirata-san researched and ordered a crane scale for use at the site. (We decided the 500-kg version of the 200-kg one used at TAMA. It is the same price and has the same 100-g resolution for the first 250 kg.)
- Hirata-san worked on making some guide parts for the PI installation (Misumi parts with minor additional machining from ATC).
- We finalized a design for simple SS mirror parts for the dummy BS with Akutsu-san. They will be made by ATC.
- Mark worked on list of items for purchase in FY2016 and on the schedule.
- Fabian did more work on the IM assembly. It is nearly done and he will start on the RM next week. Some issues with constructing pico adapters cables.
- Type B ongoing issues
- Need to decide RM wire thickness and order BS and RM wires. Lead time is a few days from Nilaco.
- Feedthrough Adapters
- Need large quantity of male and female D-Sub 9, IDC ribbon connectors - Takahashi-san is getting a quote.
- Can we make SR wire breakers bigger?
- Fukushima-san will start on machining test soon (as of 5/27).
- Not enough crane height problem (Hirata, Saito) - no action since last week
- Installation stuff
- Crane scale - ORDERED.
- Work platforms - RECEIVED.
- PI Guide parts - in manufacture at ATC.
- Ballast masses for payload
- Initial batch received - ordered extra smaller ones to allow for adjustments.
- Lower breadboard stuff
- Installation to be incorporated into the procedure document. Mark/Takahashi to discuss.
- Breadboard springs (in manufacture, expected end of July)
- Nail-head parts: quote from Mirapro received, not urgent, will be combined with assembly frame order.
- Clean bag - expected 6/8.
- Site work plan
Start the preparation work from the week of 6/20
3D CAD model (Original & Latest modified version) -> DONE: Uploaded to ROOM214 (Shoda)
1st draft of the installation procedure for modified parts -> Ongoing (Fujii+Shoda)
Rough LVDT model -> DONE: under discussion
We found that the suspension points of the optical bench are required to be shifted -> Shoda will design new adapters. (Deadline: Mid July)
Watchdog system (Tanaka)
Report of the experimental results and analysis(Resistance stability of PhotoMos Relay).http://gwdoc.icrr.u-tokyo.ac.jp/DocDB/0052/E1605270/002/Project%20WDS%20Experiment%202016_06_10.pdf
- Accelerometers for IP
NIKHEF accelerometer measurement results (Fujii) Slides
The LVDT driver had wrong power supply voltages -> need to perform the measurements again
- Type-B team will continue the noise measurements as well as the geophones
- Fujii and Shoda will go to Kamioka on 27-29 June to transfer the information to the Type-B team
- LVDT for BF damping
- If we get the design from Virgo within this month, we will use the design
- If not, we should make a design from scratch
- Interference of the interferometer beams with VIS structure
- Shoda is working on the new fixture design for PR: by mid-July
Need to check if there is a similar problem in SR (Mark)
Site work plan for the next week
BS crane will not be available: 13th -14th of June
- Travel (Week of 6/13)
- Travel (Week of 6/20):
Takahashi:19-21 (Acceptance of SFs, check items, move things around and so on)
Mark:20-24 (preparation for BS installation)
Fabian:20-24 (preparation for BS installation)
Hirata:20-24(not yet decided)
June 27-29: Fujii, Shoda
July 4 - 8: Takahashi, Ishizaki, Sato (site check for the installation of Type-A)
Mark is away: July 8-20
2016/6/17(Fri) 13:30 -
Next VIS F2F 6/17
Next next F2F 7/8